# HEAM : Hashed Embedding Acceleration using Processing-In-Memory

### 21 Feb 2024 v2

@@ 3.1. Impact on DLRM Throughput
Compositional embedding further intensifies the memory-bound characteristics of the model by doubling the overall number of accesses to the main memory. This is because generating input vectors for the model necessitates accessing two hash tables. Figure 4 (a) presents a comparison of the execution times for embedding operations on HBM and DIMM between the original model and the model with compositional embedding applied. The results indicate that compositional embedding leads to a 25% longer execution time on HBM and a 40% longer execution time on DIMM. Although the decrease in throughput is less pronounced on HBM due to its significantly higher number of channels compared to DIMM, it is evident that employing the compositional embedding results in increased memory access. This underlines the need for higher bandwidth than what traditional memory systems can provide.
3.2. Temporal and Spatial Locality
@@
To investigate the locality characteristics in compositional embedding, we simulated cache behavior with embedding traces from criteo-kaggle-dataset (criteo_kaggle_ad_dataset, ). A 4-way set associative cache is used during the experiment. Note that MLP weights are not included during the simulation, thereby producing a high cache hit rate of embeddings.
@@
Temporal Locality. To find out the presence of temporal locality, we followed the method used in RecNMP (ke2020recnmp, ). We increased cache size from 1MB to 4MB and examined the difference in hit rate of random traces and traces generated from criteo dataset (criteo_kaggle_ad_dataset, ), (reddi2020mlperf, ) with compositional embedding applied, as depicted in Figure 4 (b). The total hit rate of compositional embedding is the summation of Q table and R table hit rate, which is much higher than that of the random traces. This result implies the temporal locality characteristics of double-hashed embeddings. Although hit rate of Q table is slightly smaller than R table, it is still much higher compared to random traces, indicating strong temporal locality of Q table embeddings which resembles to original table locality characteristics examined in prior works (ke2020recnmp, ), (kal2021space, ). This is due to the similar long-tail distribution between the original and the Q table, as illustrated in Figure 5. According to Figure 5 (a), a small subset of total embeddings takes the majority of the access. We refer to this subset as hot vectors. As the number of hot vectors is small, they are sparse across the embedding table. When quotient hashing maps consecutive embeddings to the same embedding, hot vectors tend to remain as hot vectors since the access rate of Q table embedding is identical to the access rate summation of consecutive embeddings that are mapped to that Q table embedding. Therefore, long-tail distribution remains as shown in Figure 5 (b).
@@
Figure 5. Cumulative access distribution of tables. (a) Original table (b) Q table with hash collision=8.
@@
As depicted in the Figure 4 (b), R table cache hit rate occupies fixed, high portion of the total hit rate across all the cache size. This indicates each vector inside R table presents high temporal locality. The main reason is the smaller size of R table compared to Q table, even though both tables are accessed concurrently to construct the input embedding. This concentrates R table access on a small number of embeddings, significantly increasing temporal locality. It is worth noting that R table embeddings display uniform distribution as illustrated in Figure 6, which is more intensified when hash collision is low as illustrated in the figure. Since the hot vectors are randomly widespread across the embedding table, the remainder hash function maps hot vectors randomly to the R table embeddings. Coupled with the fact that R table exhibits high temporal locality, uniform access distribution of R table embeddings indicates that all the R vectors are hot vectors.
@@
Figure 6. R table access distribution of 10 embedding tables. (a) Distribution with the collision of 4 (b) Distribution with the hash collision of 8. All embeddings have similar access distribution.
@@
Spatial Locality. The simulation on spatial locality was conducted by measuring the hit rate of 64B length vector on 1MB cache, varying the cache line size from 64B to 512B on 1MB cache, again following the method used in (ke2020recnmp, ). Results are shown in Figure 4 (c). The observation shows that Q table cache hit rate decreases slightly, indicating little spatial locality. However, the R table cache hit rate implies that a mild spatial locality exists since the overall cache hit rate is stable, even though increasing cache line size diminishes the effect of temporal locality, suggesting the existence of spatial locality. Despite the presence of R table spatial locality, it is challenging to benefit from such characteristics as its impact is minimal. Furthermore, the embedding dimension exceeds traditional cache line size, as the vector length utilized in the recommendation system typically ranges from 64B to 512B. Fetching multiple R table vectors within a cache line is impossible, although R table vectors might be placed next to each other.
3.3. Opportunities for PIM Support
@@
The above assessments suggest that enhanced throughput support is necessary for running compositional embedding during the inference stage. Inspired by the previous works that utilize NMP/PIM to extend the bandwidth (kwon2019tensordimm, ), (ke2020recnmp, ), (park2021trim, ), we conclude that employing PIM is an effective solution for addressing the memory-bound characteristic of compositional embedding. Two attributes of compositional embedding make the usage of PIM more effective. First, as described in Section 2.2, the data arrangement is suitable enough to employ the PIM architecture. This is due to the uniformity in vector dimensions, which ensures efficient use of memory capacity and consistency in the total DRAM burst needed for each vector. Second, unique features of the R table vectors, high temporal locality, and its uniform memory access distribution could be efficiently exploited by extending the PIM design to increase the throughput further. Integrated with a heterogeneous system consisting of DIMM and HBM with base die PIM, the utilization of PIM is an effective design for supporting compositional embedding.
4. HEAM Architecture
Figure 7. HEAM architecture. (a) Overview of HEAM architecture, (b) 2-level PIM system in HBM
4.1. Architecture Overview
@@
We propose HEAM, a specialized memory system designed to enhance the bandwidth of DLRM when compositional embedding is employed. The overall architecture is depicted in Figure 7. As illustrated in Figure 7 (a), the HEAM system incorporates a heterogeneous memory architecture that comprises HBM and DIMM, which addresses the memory-bound nature to some extent by offloading frequently accessed vectors to HBM. Nonetheless, it’s important to note that the overall memory access demands double when employing compositional embedding. As a result, depending solely on HBM remains insufficient to achieve the required throughput.
@@
To tackle this problem, we design a two-level PIM system within the HBM to enhace the throughput by utilizing in-memory parallelism, thereby providing additional bandwidth support for the embedding operation. In Figure 7 (b), the first-level processing unit is located on the base die of the HBM, whereas the second-level processing units are located inside each bank group. These processing units are referred to as bd-PIM and bg-PIM, named based on their respective locations. Within these PIMs, the embedding operation is processed in a two-stage partial addition fashion. Initially, each bg-PIM collects Q embeddings and their corresponding R embeddings from its dedicated bank group. Local addition is performed inside each bg-PIM on the collected embedding vectors, generating partial results. Once all bg-PIMs complete their respective operations, all the reduced vectors are forwarded to the bd-PIM. The same operation is carried out within the bd-PIM for each set of partial results, resulting in the final output delivered to the host processor.
@@
Figure 8 (b) depicts the PIM architecture of bg-PIM. The system inside bg-PIM is equipped with an instruction register, an instruction decoder and a Multiply-and-Accumulate (MAC) unit. Additionally, LUT is integrated inside each bg-PIM to leverage the temporal locality of the R table described in Section 3.2. Therefore, the access to R embedding vectors is always directed to the LUT, decreasing the total bank access. HEAM employs batching of 4 embedding operations to address the load imbalance issue, following the technique utilized in RecNMP (ke2020recnmp, ). To avoid introducing too much area overhead into bank groups, we use up to 4 batch in our system. The bd-PIM receives partial sums delivered from bg-PIM and processes the final GnR operation as illustrated in Figure 8 (a). Note that the architecture of the bd-PIM is similar to that of the previous works.
Figure 8. PIM architectures. (a) bd-PIM, (b) bg-PIM
4.2. Hash Table Allocation Strategy
@@
To mitigate the double memory access problem associated with compositional embedding, we store the R and Q tables in memory in different ways. This is based on the data access patterns for each table are distinct and exhibit significant variation. Consequently, we ensure more efficient use of memory system resources. As these data access patterns are established after the training phase, we collect this distribution information prior to entering the inference stage. This additional process incurs minimal overhead, as demonstrated in previous studies such as (ke2020recnmp, ) and (kal2021space, ).
@@
Mapping Scheme for Q table. As the Q table typically exhibits a long-tail distribution similar to that of the original embedding table, we leverage an approach inspired by SPACE (kal2021space, ) to exploit the Q table’s characteristics. Embedding data with high temporal locality is offloaded to HBM, while the remaining embedding data are stored in DIMMs. The proportion of data to be offloaded to HBM is determined based on the bandwidth ratio between HBM and DIMM. This strategy, which involves gathering embedding data with high locality from HBM and fetching the rest from DIMMs, enhances the overall parallelism between the different memory types.
@@
Mapping Scheme for R table. As described in Section 3.2, access to the R table is repeatedly conducted for a very limited set of data, consequently, all the data within the R table exhibits extremely high temporal locality. Based on the observation, HEAM stores all the R table embeddings inside the LUT that is located in each bg-PIM. R embeddings are directed to the LUT inside the bg-PIM as depicted in Figure 8 (b). As a result, the number of bank accesses in compositional embedding is reduced by half, making the total number of accesses equivalent to that of the original model. Additionally, given that the R embeddings follow a uniform distribution, load balance between LUT entries is achieved automatically. The total area overhead of LUT is minimal, as the number of entries required within the LUT is typically small considering the hash collision is usually set as a value less than 100. As we employ simple addition for input embedding reconstruction operation, we evenly allocate all the R tables across all LUTs in a random fashion without forcing the Q table and R table to be placed within the same bank group.
[…]
4.3. Hash Collision Value for HEAM System
As the Q table exhibits similar locality characteristics to the original, the advantage of heterogeneous memory systems described in SPACE (kal2021space, ) is applicable to compositional embedding. Moreover, to diminish the hash collision value, the HBM-only scenario incurs redundant bandwidth since it needs to introduce additional stacks to increase capacity, while the heterogeneous system achieves the same hash collision without extra bandwidth. In order to fully reap the maximum bandwidth, each memory request ratio has to be equal to the bandwidth ratio, as displayed in the equation 1.
(1) $\frac{Request_{HBM}}{Request_{DIMM}}=\frac{BW_{HBM}}{BW_{DIMM}}$
@@ 4.3. Hash Collision Value for HEAM System
When HBM bandwidth is expanded with PIM, the overall HBM request should be enlarged to meet the above equation, forcing additional data offloading to the HBM. This might demand higher hash collision, leading to decreased model quality. However, long tail distribution compensates such problems as increasing cumulative requests of HBM requires a small amount of additional data. Therefore, there is no need to introduce extra hash collision to meet the bandwidth demand.
4.4. PIM Instructions and the Execution Flow
@@
PIM Instructions. HEAM leverages the PIM instruction to deliver embedding operation information to each bg-PIM, following RecNMP (ke2020recnmp, ) and TRiM (park2021trim, ). The total width of a single PIM instruction is 84 bits. The opcode (3-bit) specifies the operation type that bd-PIM and bg-PIM should take into action for GnR operation. The target address (34-bit) denotes the start address of the PIM-Inst to read from the bank group. nRD (4-bit) indicates the vector size. The weight (32-bit) delivers weight information for weighted sum. transfer (1-bit) determines whether the partial sum is complete and ready to deliver. The skewed-cycle (6-bit) instructs bg-PIM the time that PIM instruction should be decoded after being delivered to the bg-PIM. The batchTag (2-bit) is the assigned batch of the PIM instruction.
@@
Execution Flow. To run the PIM system, it is essential for the memory controller to deliver PIM instructions to the memory system properly. Like prior works (ke2020recnmp, ), (park2021trim, ), (kal2021space, ), HEAM offers an embedding operation driver that extends the current memory controller. The driver receives an embedding operation from the host, constructs PIM instructions, and offloads it to the NMP at the HBM. When HBM is used as a DRAM cache inside a heterogeneous memory system, a mapping of data is required to be stored inside the memory controller to direct each of the commands to the proper memory system. To support this, embeddings are mapped with a direct mapping method. With the aid of the direct mapping method, the memory controller could send PIM instructions and DDR4 commands to each memory by adopting the method used in SPACE (kal2021space, ). To avoid the cache coherence issues, we allocate memory space dedicated to HEAM as uncachable. Additionally, to cope with the command/address (C/A) bandwidth overhead issue addressed before (park2021trim, ), we follow the method or prior works that utilize data (DQ) pins to deliver a number of instructions to bg-PIMs within a short cycle. Specifically, we utilize 14 C/A pins with 256-bit wide DQ pins to transfer 1 PIM instruction in a single cycle. Due to the abundant HBM DQ pins, HEAM successfully addresses the C/A bandwidth overhead.
5. Experimental Setup

### 14 Mar 2024 v3

@@ III-A Impact on DLRM Throughput
Compositional embedding further intensifies the memory-bound characteristics of the model by doubling the overall number of accesses to the main memory. This is because generating input vectors for the model necessitates access to two hash tables. Figure 4 (a) presents a comparison of the execution times for embedding operations on HBM and DIMM between the original model and the model with compositional embedding applied. The results indicate that compositional embedding leads to a 25% longer execution time on HBM and a 40% longer execution time on DIMM. Although the decrease in throughput is less pronounced on HBM due to its significantly higher number of channels compared to DIMM, it is evident that employing the compositional embedding results in increased memory access. This underlines the need for higher bandwidth than what traditional memory systems can provide.
III-B Temporal and Spatial Locality
@@
To investigate the locality characteristics in compositional embedding, we simulated cache behavior with embedding traces from criteo-kaggle-dataset [25]. A 4-way set associative cache is used during the experiment. Note that MLP weights are not included during the simulation, thereby producing a high cache hit rate of embeddings.
@@
Temporal Locality. To find out the presence of temporal locality, we followed the method used in RecNMP [7]. We increased cache size from 1MB to 8MB and examined the difference in hit rate of random traces and traces generated from criteo dataset [25], [26] with compositional embedding applied, as depicted in Figure 4 (b). The total hit rate of compositional embedding is the summation of Q table and R table hit rate, which is much higher than that of the random traces. This result implies the temporal locality characteristics of double-hashed embeddings. Although hit rate of Q table is slightly smaller than R table, it is still much higher compared to random traces, indicating strong temporal locality of Q table embeddings, resembling original table locality characteristics that is examined in prior works [7], [9]. This is due to the similar long-tail distribution between the original and the Q table, as illustrated in Figure 5. According to Figure 5 (a), a small subset of total embeddings takes the majority of the access. We refer to this subset as hot vectors. When quotient hashing maps consecutive embeddings to the same embedding, hot vectors tend to remain as hot vectors since the access rate of Q table embedding is identical to the access rate summation of consecutive embeddings that are mapped to that Q table embedding. Therefore, long-tail distribution remains similar as shown in Figure 5 (b).
@@
Figure 5: Cumulative access distribution of tables. (a) Original table (b) Q table with hash collision=8.
@@
As depicted in the Figure 4 (b), R table cache hit rate occupies fixed, high portion of the total hit rate across all the cache size. This indicates each vector inside R table presents high temporal locality. The main reason is that R table size is smaller than that of the Q table, even though both tables are accessed concurrently to construct the input embedding. This concentrates R table access on a small number of embeddings, significantly increasing temporal locality. It is worth noting that R table embeddings display uniform locality distribution as illustrated in Figure 6, which is more intensified when hash collision is low. Since the hot vectors are randomly widespread across the embedding table, the remainder hash function maps hot vectors randomly to the R table embeddings. Coupled with the fact that total locality of the R table is high, uniform access distribution of R table embeddings indicates that all the R vectors are hot vectors.
@@
Figure 6: R table access distribution of 10 embedding tables. (a) Distribution with the collision of 4 (b) Distribution with the hash collision of 8. All embeddings have similar access distribution.
@@
Spatial Locality. The simulation on spatial locality was conducted by measuring the hit rate of 64B length vector on 1MB cache, varying the cache line size from 64B to 512B on 1MB cache, again following the method used in [7]. Results are shown in Figure 4 (c). The observation shows that Q table cache hit rate decreases slightly, indicating little spatial locality. However, the R table cache hit rate implies that a mild spatial locality exists since the overall cache hit rate is stable, even though increasing cache line size diminishes the effect of temporal locality, suggesting the existence of spatial locality. Despite the presence of R table spatial locality, it is challenging to benefit from such characteristics as its impact is minimal. Furthermore, the embedding dimension exceeds traditional cache line size, as the vector length utilized in the recommendation system typically ranges from 64B to 512B. Fetching multiple R table vectors within a cache line is impossible, although R table vectors might be placed next to each other.
III-C Opportunities for PIM Support
@@
The above assessments suggest that enhanced throughput support is necessary for running compositional embedding during the inference stage. Inspired by the previous works that utilize NMP/PIM to extend the bandwidth [6], [7], [8], we conclude that employing PIM is an effective solution for addressing the memory-bound characteristic of compositional embedding. Two attributes of compositional embedding make the usage of PIM more effective. First, as described in Section II-B, the data arrangement is suitable enough to employ the PIM architecture. This is due to the uniformity in vector dimensions, which ensures efficient use of memory capacity and consistency in the total DRAM burst needed for each vector. Second, unique features of the R table vectors, high temporal locality, and its uniform memory access distribution could be efficiently exploited by extending the PIM design to increase the throughput further. Integrated with a heterogeneous system consisting of HBM and DIMM with base die PIM, the utilization of PIM is an effective design for supporting compositional embedding.
IV HEAM Architecture
IV-A Architecture Overview
@@
We propose HEAM, a specialized memory system designed to enhance the bandwidth of DLRM when compositional embedding is employed. For the reconstruction operation of compositional embedding, addition is utilized in our system. The overall architecture is depicted in Figure 7. As illustrated in Figure 7 (a), the HEAM system incorporates a heterogeneous memory architecture that comprises HBM and DIMM, which addresses the memory-bound nature to some extent by offloading frequently accessed vectors to HBM. Nonetheless, it’s important to note that the overall memory access demands double when employing compositional embedding. As a result, depending on conventional HBM remains insufficient to achieve the required throughput.
@@
To tackle this problem, we design a two-level PIM system within the HBM to enhace the throughput by utilizing in-memory parallelism, thereby providing additional bandwidth support for the embedding operation. In Figure 7 (b), the first-level processing unit is located on the base die of the HBM, whereas the second-level processing units are located inside each bank group. These processing units are referred to as bd-PIM and bg-PIM, named based on their respective locations. Within these PIMs, the embedding operation is processed in a two-stage partial addition fashion. Initially, each bg-PIM collects Q embeddings and R embeddings from its dedicated bank group. Local addition is performed inside each bg-PIM on the collected embedding vectors, generating partial results. Note that Q embeddings and their corresponding R embeddings don’t have to be located in the same bank group, as addition operation is employed, resulting associative law property to be held in the operation. Once all bg-PIMs complete their respective operations, all the reduced vectors are forwarded to the bd-PIM. The same operation is carried out within the bd-PIM for each set of partial results, resulting in the final output delivered to the host processor.
Figure 7: HEAM architecture. (a) Overview of HEAM architecture, (b) 2-level PIM system in HBM
@@
Figure 8 (b) depicts the PIM architecture of bg-PIM. The system inside bg-PIM is equipped with an instruction register, an instruction decoder and a multiply-and-accumulate (MAC) unit. Additionally, LUT is integrated inside each bg-PIM to leverage the temporal locality of the R table described in Section III-B. Therefore, the access to R embedding vectors is always directed to the LUT, decreasing the total bank access. HEAM employs batching of 4 embedding operations to address the load imbalance issue, following the technique utilized in RecNMP [7]. Batching compensates load imbalance issue of the embedding lookup, and the level of compensation increases with large vector size as operation overlap cycles between bank group also increases. To avoid introducing too much area overhead into bank groups, we use up to 4 batch in our system. The bd-PIM receives partial sums delivered from bg-PIM and processes the final GnR operation as illustrated in Figure 8 (a). Note that the architecture of the bd-PIM is similar to that of the previous works.
Figure 8: PIM architectures. (a) bd-PIM, (b) bg-PIM
IV-B Hash Table Allocation Strategy
@@
To mitigate the double memory access problem associated with compositional embedding, we store the Q and R tables in memory in different ways. This is based on the data access patterns for each table are distinct and exhibit significant variation. Consequently, we ensure more efficient use of memory system resources. As these data access patterns are established after training, we collect this distribution information prior to entering the inference stage. This process incurs minimal overhead in time, as demonstrated in previous studies [7], [9].
@@
Mapping Scheme for Q table. As the Q table typically exhibits a long-tail distribution similar to that of the original embedding table, we leverage an approach inspired by SPACE [9] to exploit the Q table’s characteristics. Q embeddings with high temporal locality is offloaded to HBM, while the remaining Q embeddings are stored in DIMMs. The proportion of data to be offloaded to HBM is determined based on the bandwidth ratio between HBM and DIMM. This strategy, which involves gathering embedding data with high locality from HBM and fetching the rest from DIMMs, enhances the overall parallelism between the different memory types.
@@
Mapping Scheme for R table. As described in Section III-B, access to the R table is repeatedly conducted for a very limited set of data, consequently, all the data within the R table exhibits extremely high temporal locality. Based on the observation, HEAM stores all the R table embeddings inside the LUT that is located in each bg-PIM. Access to R embeddings are directed to the LUT inside the bg-PIM as depicted in Figure 8 (b). As a result, the number of bank accesses in compositional embedding is reduced by half, making the total number of accesses equivalent to that of the original model. Additionally, given that the R embeddings follow a uniform distribution, load balance between LUT entries is achieved automatically. The total area overhead of LUT is minimal, as the number of entries required within the LUT is typically small considering the hash collision is usually set as a value less than 60 [19]. As we employ simple addition for input embedding reconstruction operation, we evenly allocate all the R tables across all LUTs in a random fashion without forcing the Q table and R table to be placed within the same bank group.
While bd-PIM could also be considered for LUT placement, bg-PIM offers distinct advantages for extending reconstruction operations for multiplication and concatenation case, as those cases require q vector and its corresponding r vector to be located within the same bank group since associtave law doesn’t hold in those operations.
IV-C Advantage of HEAM System in Hash Collision Value
In the HBM-only scenario, when the model size is larger than HBM capacity, additional stacks should be introduced to avoid high hash collision value that might lead to excessive model quality drop of the model. However, this comes at the cost of redundant bandwidth. Employing an HBM-DIMM heterogeneous system is a better approach as it achieves the same hash collision without extra redundant bandwidth. To fully reap the maximum bandwidth of a heterogeneous memory system, each memory request ratio has to be equal to the bandwidth ratio, as displayed in the equation 1.
$\frac{Request_{HBM}}{Request_{DIMM}}=\frac{BW_{HBM}}{BW_{DIMM}}$ (1)
@@ IV-C Advantage of HEAM System in Hash Collision Value
However, when HBM bandwidth is expanded with PIM, the overall HBM request should be enlarged to meet the above equation, which forces additional data offloading to the HBM. If total data is increased more than HBM capacity, this might demand higher hash collision, leading to decreased model quality. Nevertheless, long tail distribution compensates for such problems as increasing cumulative requests of HBM requires a small amount of additional data. Therefore, there is no need to introduce extra hash collision to meet the bandwidth demand.
IV-D PIM Instructions and the Execution Flow
@@
PIM Instructions. HEAM leverages the PIM instruction to deliver embedding operation information to each bg-PIM, following RecNMP [7] and TRiM [8]. The total width of a single PIM instruction is 81 bits. The opcode (3-bit) specifies the operation type that bd-PIM and bg-PIM should take into action for GnR operation. The target address (33-bit) denotes the start address of the PIM-Inst to read from the bank group. nRD (4-bit) indicates the vector size. The weight (32-bit) delivers weight information for weighted sum. transfer (1-bit) determines whether the partial sum is complete and ready to deliver. The skewed-cycle (6-bit) instructs bg-PIM the time that PIM instruction should be decoded after being delivered to the bg-PIM. The batchTag (2-bit) is the assigned batch of the PIM instruction.
@@
Execution Flow. To run the PIM system, it is essential for the memory controller to deliver PIM instructions to the memory system properly. Like prior works [7], [8], [9], HEAM offers an embedding operation driver that extends the current memory controller. The driver receives an embedding operation from the host, constructs PIM instructions, and offloads it to the NMP at the HBM. When HBM is used as a DRAM cache inside a heterogeneous memory system, a mapping of data is required to be stored inside the memory controller to direct each of the commands to the proper memory system. To support this, embeddings are mapped with a direct mapping method. With the aid of the direct mapping method, the memory controller could send PIM instructions and DDR4 commands to each memory by leveraging the method used in SPACE [9]. To avoid the cache coherence issues, we allocate memory space dedicated to HEAM as uncachable. Additionally, to cope with the command/address (C/A) bandwidth overhead issue addressed before [8], we follow the method of prior works that utilize data (DQ) pins to deliver a number of instructions to bg-PIMs within a short cycle. Specifically, we utilize 14 C/A pins with 128-bit wide DQ pins to transfer 1 PIM instruction in a single cycle. Due to the abundant HBM DQ pins, HEAM successfully addresses the C/A bandwidth overhead.
V Experimental Setup